OpenStudy (anonymous):

Simplify the following 2sin15(cos15)

OpenStudy (anonymous):

@ranga @hartnn

hartnn (hartnn):

use the double angle formula \(2\sin x\cos x = \sin 2x \)

OpenStudy (anonymous):

But what values nothing for 15 degrees not like our 30 or 60 values ?

hartnn (hartnn):

no, 15 degree is not a standard angle

hartnn (hartnn):

\(2\sin15(\cos15) = \sin (2 \times 15) =...\)

OpenStudy (anonymous):

30

OpenStudy (anonymous):

sin 2(15)=0.5

hartnn (hartnn):

yes and sin 30 is indeed 0.5 :)

OpenStudy (anonymous):

\[\frac{1}{2}\]

hartnn (hartnn):

any doubts ?

OpenStudy (anonymous):

no,sir